### v3

# Towards Generalizable Scene Change Detection

Abstract
While current state-of-the-art Scene Change Detection (SCD) approaches achieve impressive results in well-trained research data, they become unreliable under unseen environments and different temporal conditions; in-domain performance drops from 77.6% to 8.0% in a previously unseen environment and to 4.6% under a different temporal condition—calling for generalizable SCD and benchmark. In this work, we propose the Generalizable Scene Change Detection Framework (GeSCF), which addresses unseen domain performance and temporal consistency—to meet the growing demand for anything SCD. Our method leverages the pre-trained Segment Anything Model (SAM) in a zero-shot manner. For this, we design Initial Pseudo-mask Generation and Geometric-Semantic Mask Matching—seamlessly turning user-guided prompt and single-image based segmentation into scene change detection for a pair of inputs without guidance. Furthermore, we define the Generalizable Scene Change Detection (GeSCD) benchmark along with novel metrics and an evaluation protocol to facilitate SCD research in generalizability. In the process, we introduce the ChangeVPR dataset, a collection of challenging image pairs with diverse environmental scenarios—including urban, suburban, and rural settings. Extensive experiments across various datasets demonstrate that GeSCF achieves an average performance gain of 19.2% on existing SCD datasets and 30.0% on the ChangeVPR dataset, nearly doubling the prior art performance. We believe our work can lay a solid foundation for robust and generalizable SCD research.
Figure 1: Comparative results of the current state-of-the-art model and our GeSCF under various unseen environments (ChangeVPR). GeSCF outperforms with precise boundaries and edges, where the state-of-the-art model hardly captures changes.
* Corresponding author
1 Introduction
Scene Change Detection (SCD) [37] is a pivotal technology that enables a wide range of applications: visual surveillance [54], anomaly detection [20], mobile robotics [31], and autonomous vehicles [21]. The ability to accurately identify meaningful changes in a scene across different time steps—despite challenges such as illumination variations, seasonal changes, and weather conditions—plays a key role in the system’s effectiveness and reliability.
Recent SCD models have achieved remarkable improvement by leveraging deep features [3, 48] and advancing model architectures [44, 11, 35, 50]. However, this progress raises a fundamental question: “Can these models detect arbitrary real-world changes beyond the scope of research data?” Our findings, as shown in Fig. 1, indicate that their reported effectiveness does not hold in real-world applications. Specifically, we observe that they produce inconsistent change masks when the input order is reversed and exhibit significant performance drops when deployed to unseen domains with different visual features. It is because current SCD methods rely heavily on their training datasets, which are often limited in size [3, 42], sparse in coverage [44, 3, 42], and predominantly synthetic [34, 35, 25] due to the costly change annotation.
To address these challenges, we introduce the Generalizable Scene Change Detection Framework (GeSCF), the first zero-shot scene change detection method that operates robustly regardless of the temporal input order and environmental conditions. Our approach builds upon the Segment Anything Model (SAM) [23], a pioneering vision foundation model for image segmentation. While SAM excels at segmenting anything within a single image, directing SAM to identify and segment changes between two input images presents a significant challenge. This difficulty arises because SAM is designed for promptable interactive segmentation, relying on user-guided prompts and single-image inputs, whereas scene change detection necessitates processing image pairs to identify changes. To bridge this gap, we propose two innovations: the Initial Pseudo-mask Generation and the Geometric-Semantic Mask Matching. By analyzing the localized semantics of the SAM’s feature space, we effectively binarize pixel-level change candidates with zero additional cost. Additionally, we leverage the geometric properties of SAM’s class-agnostic masks and the semantics of mask embeddings to refine the final change masks—incorporating object-level information as well.
Furthermore, we introduce the Generalizable Scene Change Detection (GeSCD) benchmark by developing novel metrics and an evaluation protocol to foster SCD research in generalizability; most conventional SCD methods have focused on individual benchmarks separately rather than the generalizability to unseen domains and temporal consistency. We believe our GeSCD can meet the growing needs for developing anything SCD in the era of diverse anything models [23, 66, 58, 51, 61, 56] with strong zero-shot capabilities. Concretely, our GeSCD performs extensive cross-domain evaluations to rigorously test the generalizability across diverse environments and assesses temporal consistency quantitatively. Our dual-focused evaluation strategy not only ensures the robustness and reliability of a method but also sets a new benchmark in the SCD field. In the process of designing GeSCD, we collect the ChangeVPR dataset, which comprises carefully annotated images sourced from three prominent Visual Place Recognition (VPR) datasets. This comprehensive dataset includes urban, suburban, and rural environments under challenging conditions, significantly expanding the traditional scope of SCD domains.
To summarize, our contributions are as follows:
  1. 1.
    Problem Formulation. We introduce GeSCD, a novel task formulation in scene change detection. To the best of our knowledge, this is the first comprehensive effort to address the generalization problem and temporal consistency in SCD research.
  2. 2.
    Model Design. To tackle the GeSCD task, we propose GeSCF, the first zero-shot scene change detection model. GeSCF exhibits complete temporal consistency and demonstrates strong generalizability over previous SCD models that are tightly coupled to their training datasets—ours resulting in a substantial performance gain in unseen domains.
  3. 3.
    Benchmark Set up. We present new evaluation metrics, the ChangeVPR dataset, and an evaluation protocol that effectively measures an SCD model’s generalizability. These contributions provide a solid foundation to guide and inspire future research in the field.
2 Related Works
Segment Anything Model. Segment Anything Model (SAM) [23] has set a new standard in image segmentation and made significant strides in various computer vision domains: medical imaging [63], camouflaged object detection [46], salient object detection [27], image restoration [61], image editing [56], and video object tracking [58]. By leveraging geometric prompts—such as points or bounding boxes—SAM showcases exceptional zero-shot transfer capabilities across diverse segmentation tasks and unseen image distributions. While SAM demonstrates impressive abilities, its potential for zero-shot scene change detection remains largely unexplored. In this work, we extend SAM’s utility beyond single-image segmentation by introducing a novel, training-free approach that guides SAM to detect changes between a pair of natural images.
Change Detection. In the field of Change Detection (CD), the research falls broadly into three areas based on data characteristics: remote sensing CD, video sequence CD, and natural scene CD—the focus of our work. Remote sensing CD [41, 10, 7, 19, 5, 6, 32] involves detecting surface changes over time using data captured by satellite or aerial platforms, providing a high-altitude perspective of phenomena such as urbanization, deforestation, and disaster damage. Additionally, video sequence CD focuses on segmenting frames into foreground and background regions, usually corresponding to moving objects [28, 2]. In contrast to these CDs, natural scene CD aims to detect localized changes from a ground-level perspective such as movement of vehicles [3], pedestrians [44], the appearance and disappearance of objects [3, 44, 42], and significant background changes like the construction or demolition of buildings [42, 44]. Moreover, the task inherently involves misaligned and noisy images due to the nature of data acquisition, as images are often captured by cameras mounted on moving vehicles or robots [26]. Overall, our work concentrates natural scene CD [50, 35, 3, 11, 43, 44, 48, 25]; for the remainder of this paper, we will refer to natural Scene CD as SCD.
Scene Change Detection (SCD). In exsiting SCD benchmarks, most methods are supervised [44, 11, 50, 35, 3, 43, 48] or semi-supervised [25], heavily optimized and evaluated on specific training datasets—leading to low generalizability. Although several works have proposed self-supervised pre-training strategies [39] or leveraged temporal symmetry [50], they still exhibit significant performance gaps when deployed to unseen data. Moreover, the symmetric structure relies on a specific prior knowledge of the domain, rendering it impractical for unknown domains without the proper inductive bias. In contrast, our GeSCF stands as a unified, training-free framework displaying robust performance on unseen data while preserving symmetric architecture for all settings.
Segment Anything with Change Detection. Previous works have primarily leveraged SAM in the remote sensing CD with the Parameter-Efficient Fine-Tuning (PEFT) strategy [57]. For instance, several works leveraged SAM variants [64, 62] with learnable adaptors [57], fine-tuning adaptor networks and change decoders tailored to specific datasets [13, 29]. In contrast to these approaches, our method is the first solid SAM-integrated framework in SCD, leveraging SAM’s internal byproducts to effectively binarize change candidates without any guidance and learnable parameters. Moreover, we further exploit valuable prior information [61, 56, 1, 18, 12] from SAM’s class-agnostic masks—enabling robust zero-shot SCD across a broad range of domains for the first time.
Figure 2: Illustration of the proposed GeSCF pipeline. The GeSCF pipeline consists of two major steps: (1) initial pseudo-mask generation and (2) geometric-semantic mask matching. First, we intercept facet features from the SAM image encoder and correlate them to obtain multi-head similarity maps, which are then converted into pseudo-masks using an adaptive threshold function. Next, SAM’s class-agnostic masks and last image embeddings are utilized to refine these pseudo-masks based on both geometric and semantic information.
3 GeSCF
3.1 Motivation and Overview
Despite the abundance of web-scale data [36, 33, 8] and the advent of various zero-shot generalizable models [23, 60, 22], current SCD still suffers from the curse of datasets due to the costly nature of change annotation [65]. Therefore, our research motivation arises from how SCD can benefit from recent web-scale trained models like SAM. By addressing this question, we aim to overcome the longstanding obstacle of creating a generalizable SCD model—culminating in our proposed GeSCF model.
Fig. 2 gives an overview of the GeSCF pipeline. GeSCF handles the technical gap between SAM designed for promptable interactive segmentation with single-image inputs and SCD for identifying changes with image pairs through two key stages: Initial Pseudo-mask Generation and Geometric-Semantic Mask Matching. First, we intercept and correlate feature facets—one of query, key, and value—from the image encoder to obtain rich, multi-head similarity maps; then, we transform these similarity maps into a binary pseudo-mask by adaptively thresholding the low-similarity pixels using a skewness-based algorithm. Finally, we elaborate the pseudo-mask by leveraging the geometric properties of SAM’s class-agnostic masks; then, we further validate these masks by comparing the semantic similarities of the corresponding mask embeddings between bi-temporal images, ensuring that the detected changes are meaningful and contextually accurate.
3.2 Preliminary
Since our GeSCF exploits a set of image features intercepted from the SAM image encoder, we first recap how such features are obtained.
Feature Facets. The SAM image encoder employs a Vision Transformer (ViT) architecture [15], consisting of multi-head self-attention layers and multi-layer perceptrons within each ViT block [15, 49]. In the multi-head self-attention layer of the $l$-th ViT block, the query, key, and value facets are denoted by $\mathbf{QKV}_{l}\in\mathbb{R}^{3\times N\times H\times W\times C}$, where $N$, $H$, $W$, and $C$ represent the number of heads, height, width, and channel dimensions of facets, respectively.
Image Embedding and Mask Embedding. Similarly, we extract the image embedding $\mathbf{E}_{l}\in\mathbb{R}^{H\times W\times C}$ from the final multi-layer perceptron layer of the $l$-th ViT block. Furthermore, given the image embedding $\mathbf{E}_{l}$ and an arbitrary binary mask $\overline{\mathbf{m}}$, we compute the mask embedding $\mathcal{M}_{l}$ by averaging image embedding $\mathbf{E}_{l}$ over all spatial positions where the binary mask $\overline{\mathbf{m}}$ is non-zero, obtaining a single vector representation of the masked image.
3.3 Initial Pseudo-mask Generation
As demonstrated in [23], SAM preserves semantic similarities among mask embeddings within the same natural scene. Furthermore, as observed in prior studies [4, 9], attention maps can capture semantically meaningful objects in images. Building on these foundational insights, we expand SAM’s feature space utilization by extending its application to bi-temporal images and leveraging multi-head feature facets from various layers rather than relying solely on single-image embeddings [23].
Multi-head Feature Correlation. Given a bi-temporal RGB image pair, we first estimate a coarse transformation [26, 47, 24] relating the bi-temporal images. Then, we intercept the feature facets $\mathbf{F}_{l,n}^{t_{0}}$, $\mathbf{F}_{l,n}^{t_{1}}$ (one of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$) for the $n$-th head from the $l$-th ViT block and compute multi-head feature correlation as follows:
$\overline{\mathbf{S}}_{l,n}^{t_{0}\leftrightarrow t_{1}}=\mathbf{F}_{l,n}^{t_{0}}\prescript{\top}{}{\mathbf{F}_{l,n}^{t_{1}}},$ (1)
$\mathbf{S}_{l}^{t_{0}\leftrightarrow t_{1}}=\xi\left(\sum_{n=1}^{N}\overline{\mathbf{S}}_{l,n}^{t_{0}\leftrightarrow t_{1}}\right)\vspace{0.1cm},$ (2)
where $\xi(\cdot)$ performs spatial reshaping and bilinear interpolation to the input image size. The correlation in Eq. 1 is a commutative operation that generates identical similarity maps $\overline{\mathbf{S}}_{l,n}$ even with the reversed order. Subsequently, the similarity maps are averaged over the head dimension—leveraging the semantic diversity across the $N$ heads.
Facet and Layer Selection. As shown in Fig. 3 , the similarity map of SAM’s facets effectively highlights semantic changes while remaining relatively unaffected by visual variations, such as seasonal or illumination differences. Concretely, our feature selection strategy is guided by the following principles: (a) semantic changes should appear prominently in the similarity map compared to other visual variations; (b) the similarity values in changed regions should clearly contrast with surrounding areas; and (c) artifacts in unchanged regions should be minimized or appear faint. We empirically observe that all facets satisfy principle (a); however, principles (b) and (c) are particularly noticeable when using the key (or query) facet over the value facet. Moreover, we find that these principles hold more strongly in intermediate layers than in the initial or last layers. Consequently, we leverage the similarity map of key facets from intermediate layers as inputs to the subsequent steps in the pseudo-mask generation. For the detailed quantitative analysis, please refer to the supplementary material.
Figure 3: Visualization of the similarity map depending on the facets and the layers. We use key facets from the intermediate layer of the SAM ViT image encoder, which is highlighted with a red bounding box.
Adaptive Threshold Function. To create a binary mask from an arbitrary similarity map, the most straightforward approach is to apply a fixed threshold (e.g., $p=0.5$) [17], where similarity scores below this threshold are classified as changes. However, this fixed threshold approach is inherently limited, as it fails to account for the relative nature of “change” within each similarity map. For instance, a similarity score of 0.7 may signify a change if all other values are close to one, whereas it may not indicate a change if most values are near zero. Consequently, the perception of “change” is context-dependent, necessitating an adaptive thresholding method that accounts for the relative distribution of similarity scores within each map.
Therefore, a critical factor to binarize the similarity map is the skewness ($\gamma$) [38] of the similarity distribution (see Fig. 4). For right-skewed distributions, where the majority of pixels display lower similarity scores with a long tail of higher scores, a lower threshold is required to capture a larger portion of the distribution as changes. Conversely, for left-skewed distributions, where most scores are high with a small tail of lower values, a higher threshold is needed to avoid false positives. To this end, we propose an adaptive threshold function for dynamic adjustment based on the skewness of the distribution—enabling a more accurate and context-sensitive method for pseudo-mask generation, as detailed in the following formulation:
$\mathbf{F}(\gamma)=b_{\gamma}+c\cdot\operatorname{sign}(\gamma)\cdot s_{\gamma}\cdot\gamma,$ (3)
where $b_{\gamma}$ is the baseline threshold, $c$ is a normalization constant, $s_{\gamma}$ is the skewness sensitivity factor, $\gamma$ is the skewness of the distribution, and $\operatorname{sign}(\cdot)$ adjusts the threshold direction based on the skewness. By applying the computed adaptive threshold to the similarity map—normalized using the mean absolute deviation (MAD) [38]—we obtain the initial pseudo-mask essential for the subsequent Geometric-Semantic Mask Matching.
Figure 4: Illustration of the adaptive thresholding process. We dynamically adjust the threshold based on the skewness of the distribution to generate the initial pseudo-masks.
3.4 Geometric-Semantic Mask Matching
Building upon the initial pseudo-masks, we elevate our focus to detecting object-level changes by using SAM’s class-agnostic object proposals. This transition from pixel-wise analysis to object-level consideration enables a more comprehensive and interpretable change mask. Here, we introduce two matching strategies: geometric intersection matching (GIM) and semantic similarity matching (SSM).
Geometric Intersection Matching. The main idea of this strategy is to select SAM masks by evaluating their overlap with the initial pseudo-masks. We calculate the intersection ratio $\alpha$ between each SAM mask and the pseudo-mask, retaining only those masks with $\alpha$$>$$\alpha_{t}$, where $\alpha_{t}$ is a threshold. Since changes can occur bi-temporally, we apply the process for both $t_{0}$ and $t_{1}$ images—maintaining commutativity for the proposed GeSCF.
Semantic Similarity Matching. Although GIM provides reasonable masks for potential object-level changes, a set of unchanged masks is included due to noise in the initial pseudo-masks. To address this issue, we perform semantic verification of the overlapping regions by calculating the cosine similarity between the mask embeddings of the bi-temporal images. Specifically, for each overlapping mask $\overline{m}_{o}$, we extract the corresponding mask embeddings $\mathcal{M}_{l,o}^{t_{0}}$ and $\mathcal{M}_{l,o}^{t_{1}}$ from the $l$-th image embeddings $\mathbf{E}_{l}$ at times $t_{0}$ and $t_{1}$, respectively. We then compute a change confidence score using the cosine similarity $c(\mathcal{M}_{l,o}^{t_{0}},\mathcal{M}_{l,o}^{t_{1}})$ which allows further refinement of the masks selected by GIM and generates the final change mask $\mathbf{Y}_{\text{pred}}$. Through a layer-wise analysis, we empirically observe that semantic differences are more pronounced in the last layer compared to the initial and intermediate layers, thereby utilizing final image embeddings for our SSM process [^1] Please refer to the supplementary material for the analysis..
4 GeSCD
Since the introduction of the first SCD benchmark [42] in 2015, the generalizability to unseen domains and temporal consistency of predictions have not been consistently or comprehensively addressed in the SCD field. Most traditional SCD approaches train and evaluate models on individual datasets separately [50, 35, 3, 11, 43, 44, 48, 25]. Only [39] performed cross-domain evaluation, however, it still suffers from domain gaps and limited training datasets. Moreover, in contrast to the remote sensing CD [65], the temporal consistency is often overlooked in model design [35, 3, 11, 43, 44, 48, 25] or training objectives [50, 35, 3, 11, 43, 44, 48, 25] in the SCD field. The temporal symmetry proposed in [50] is unsuitable for real-world applications since it assumes perfect inductive bias of the application domain, which is not feasible in real-world scenarios. Further, as we are in the age of diverse anything models [23, 66, 58, 51, 61, 56] with strong zero-shot prediction and generalizability, the necessity of anything SCD model that can adapt to various change scenarios has become increasingly important in the research community.
Based on these requirements, we propose GeSCD—a novel task approach that addresses the generalizability of broader scenarios and temporal consistency of SCD models. By pioneering new metrics, an evaluation dataset, and a comprehensive evaluation protocol, our approach fulfills the critical need for SCD research that is genuinely applicable and effective across diverse settings.
[…]
where the proposed TC score indicates how much the SCD algorithms can generate consistent change masks in bi-directional orders.
Evaluation Datasets. First, we consider three standard SCD datasets with different characteristics: VL-CMU-CD [3], TSUNAMI [42], and ChangeSim [34]. These datasets represent urban environments in the USA, disaster-impacted urban areas in Japan, and industrial indoor settings within simulation environments, respectively. Furthermore, for the quantitative evaluation on broader unseen domains, we create a new dataset named the ChangeVPR. The ChangeVPR comprises 529 image pairs collected from the SF-XL (urban, U) [8], St Lucia (suburban, S) [30], and Nordland (rural, R) [45] datasets, which are widely used in the Visual Place Recognition (VPR) research. We carefully sampled image pairs from each dataset to reflect various SCD challenges such as weather conditions and seasonal changes, and hand-labeled the ground-truth change masks for each image pair. We employ ChangeVPR only for the evaluation to assess unseen domain performance. Please refer to the supplementary material for more details.
Evaluation Protocol. We perform extensive cross-domain evaluation for SCD models. First, We train each model on the three standard SCD datasets; the training stage results in three distinctive models for each method. Then, we assess each of the three distinctive models in two stages. First, we evaluate the three models on the three conventional SCD datasets; this process totals nine assessments. If the method of interest does not involve training, we evaluate the model on the three datasets (three assessments). Next, we evaluate the models on the proposed ChangeVPR dataset. Since we use ChangeVPR only for evaluation, this stage contains three assessments per model (due to three splits of ChangeVPR). As a result, the proposed protocol estimates performance on both seen and unseen domains—ensuring a thorough evaluation of SCD performance in various environmental scenarios with arbitrary real-world changes.
5 Experiments
Method Training VL-CMU-CD TSUNAMI ChangeSim Avg.
[…]
Uni-temporal Training
CSCDNet VL-CMU-CD 77.4 4.5 0.02 5.6 19.4 0.02 25.5 13.6 0.09 24.3
CDResNet 74.7 2.1 0.01 2.3 4.6 0.0 25.4 13.2 0.12 20.4
DR-TANet 74.2 1.6 0.01 2.6 6.7 0.0 27.5 18.0 0.19 21.8
C-3PO 77.6 4.6 0.02 8.0 34.3 0.02 30.4 21.0 0.16 29.3
CSCDNet TSUNAMI 17.0 22.1 0.40 83.6 60.3 0.46 27.8 30.5 0.29 40.2
CDResNet 13.2 18.4 0.42 82.8 51.5 0.38 26.3 29.0 0.31 36.9
DR-TANet 14.8 17.3 0.53 82.0 57.6 0.44 24.4 26.8 0.35 37.2
C-3PO 27.0 27.0 1.0 84.2 84.2 1.0 34.3 34.3 1.0 48.5
CSCDNet ChangeSim 20.1 3.5 0.04 2.5 6.6 0.01 43.1 18.9 0.31 15.8
CDResNet 17.5 3.1 0.02 4.6 9.9 0.01 41.3 18.2 0.18 15.8
DR-TANet 20.3 4.9 0.05 5.0 8.3 0.01 40.3 18.0 0.18 16.1
C-3PO 1.6 1.6 1.0 0.3 0.3 1.0 15.8 15.8 1.0 5.9
Bi-temporal Training
CSCDNet VL-CMU-CD 64.2 64.8 0.61 8.5 7.5 0.12 24.0 23.9 0.42 32.2
CDResNet 60.6 61.7 0.58 5.0 3.8 0.18 17.7 14.1 0.29 27.2
DR-TANet 45.6 61.9 0.40 3.5 1.2 0.02 18.0 23.0 0.26 25.5
C-3PO 58.4 63.5 0.63 2.5 1.4 0.07 21.2 23.8 0.51 28.5
CSCDNet TSUNAMI 20.9 21.4 0.56 82.4 82.8 0.82 30.0 29.4 0.46 44.5
CDResNet 17.6 17.1 0.63 81.7 81.8 0.81 28.9 29.0 0.48 42.7
DR-TANet 18.1 17.4 0.61 81.1 82.0 0.76 28.6 28.7 0.46 42.7
C-3PO 27.5 27.5 1.0 84.1 84.1 1.0 33.7 33.7 1.0 48.4
CSCDNet ChangeSim 0.2 0.1 0.86 4.7 1.0 0.69 19.5 18.6 0.49 7.4
CDResNet 4.0 2.0 0.75 1.1 0.5 0.53 19.8 20.7 0.48 8.0
DR-TANet 7.5 3.7 0.48 2.6 1.8 0.13 27.9 21.4 0.32 10.8
C-3PO 0.4 0.4 1.0 1.2 1.2 1.0 18.0 18.0 1.0 6.5
GeSCF (Ours) Zero-shot 75.4 75.4 1.0 72.8 72.8 1.0 54.8 54.8 1.0 67.7
Table 1: Quantitative results (F1-score) on standard SCD datasets.
Method Training SF-XL (U) St Lucia (S) Nordland (R) Avg.
[…]
Uni-temporal Training
CSCDNet VL-CMU-CD 28.6 23.1 0.07 11.7 15.3 0.27 14.3 5.8 0.33 16.5
CDResNet 22.9 18.9 0.08 13.5 18.0 0.18 9.3 6.7 0.39 14.9
DR-TANet 22.5 19.6 0.05 15.5 20.8 0.19 11.8 8.2 0.33 16.4
C-3PO 38.7 32.8 0.07 21.7 25.7 0.14 16.2 14.2 0.28 24.9
CSCDNet TSUNAMI 38.9 39.6 0.55 15.3 15.4 0.55 18.8 20.8 0.35 24.8
CDResNet 36.6 37.4 0.59 13.7 14.0 0.66 16.4 19.9 0.33 23.0
DR-TANet 35.2 36.1 0.62 13.4 13.4 0.76 17.1 16.7 0.44 22.0
C-3PO 50.0 50.0 1.0 28.4 28.4 1.0 23.2 23.2 1.0 33.9
CSCDNet ChangeSim 32.8 28.1 0.08 19.3 19.6 0.09 6.7 7.8 0.29 19.1
CDResNet 26.6 23.4 0.07 15.2 14.2 0.15 7.1 6.4 0.3 15.5
DR-TANet 27.5 25.0 0.08 15.9 15.9 0.15 10.6 13.8 0.21 18.1
C-3PO 4.0 4.0 1.0 3.1 3.1 1.0 0.8 0.8 1.0 2.6
Bi-temporal Training
CSCDNet VL-CMU-CD 24.1 25.4 0.40 5.6 5.1 0.68 2.2 3.0 0.71 10.9
CDResNet 23.1 23.3 0.37 7.0 6.7 0.69 1.3 1.2 0.80 10.4
DR-TANet 17.0 17.3 0.27 9.0 4.1 0.56 2.2 2.4 0.66 8.7
C-3PO 23.3 23.2 0.5 7.9 6.9 0.77 3.6 4.6 0.79 11.6
CSCDNet TSUNAMI 42.8 42.9 0.67 17.1 17.1 0.64 20.4 21.0 0.57 26.9
CDResNet 40.0 40.1 0.71 13.9 13.8 0.76 19.8 19.5 0.58 24.5
DR-TANet 38.6 38.7 0.71 13.9 14.0 0.78 20.6 17.5 0.55 23.9
C-3PO 50.0 50.0 1.0 28.0 28.0 1.0 24.3 24.3 1.0 34.1
CSCDNet ChangeSim 13.3 12.1 0.31 0.9 1.1 0.82 0.7 1.0 0.84 4.9
CDResNet 11.4 9.6 0.26 2.2 1.5 0.76 1.3 0.7 0.92 4.5
DR-TANet 12.8 12.5 0.12 8.3 6.1 0.30 7.5 6.3 0.40 8.9
C-3PO 6.0 6.0 1.0 2.1 2.1 1.0 1.0 1.0 1.0 3.0
GeSCF (Ours) Zero-shot 71.2 71.2 1.0 62.1 62.1 1.0 59.0 59.0 1.0 64.1
Table 2: Quantitative results (F1-score) on the ChangeVPR dataset.
Figure 5: Qualitative results on VL-CMU-CD dataset with F1-scores. Our model generates reasonable change masks and does not display annotation bias.
5.1 Comparative Studies
We compare our method with four state-of-the-art SCD models: CSCDNet [44], CDResNet [44], DR-TANet [11], and C-3PO [50]. For C-3PO, we adopt the (I+D) structure for the VL-CMU-CD and the (I+A+D+E) structure for the TSUNAMI and ChangeSim datasets, following the configurations proposed in the paper. To verify the temporal consistency, we additionally train all models with bi-temporal objective [65]; the performance of these bi-temporally trained models has not been reported in the literature.
Quantitative Comparison. Tab. 1 displays the quantitative comparison results on the standard SCD datasets. The results attest that GeSCF outperforms the baselines with a large margin in every unseen domain (off-diagonal results) and on-par performance on seen domains (diagonal results), along with complete TC score (TC$=$$1.0$) for all settings—demonstrating its superior generalizability over diverse environments and temporal conditions. The current best performance on the VL-CMU-CD dataset drops sharply from $77.6\%$ to $4.6\%$ when the temporal order is reversed and to $8.0\%$ when deployed to the unseen TSUNAMI dataset. Remarkably, our GeSCF even outperforms the supervised baselines for the ChangeSim dataset by an F1-score of 54.8% in a zero-shot manner. Overall, the average performance of our method on standard SCD datasets surpasses that of the second-best by a substantial margin ($+19.2\%$). Furthermore, the majority of baselines do not display a complete TC score which is crucial for the system’s reliability. The temporal consistency of C-3PO does not hold in VL-CMU-CD settings as shown in its low TC scores. We find that simple bi-temporal training does not guarantee complete temporal consistency and unseen domain performance, even worsening the uni-temporal performances—indicating different data characteristics from [65] and the significance of our commutative architecture. Tab. 2 presents the quantitative comparison results on the proposed ChangeVPR dataset. The results consistently show that GeSCF outperforms the baselines across all unseen domains with a $+30.0\%$ margin (nearly doubling the prior art performance), extending beyond the conventional urban-/synthetic-only SCD scope. Notably, GeSCF demonstrates exceptional performance in Nordland [45] split, a challenging dataset with severe seasonal variation (summer-winter).
Qualitative Comparison. As illustrated in Fig. 1, GeSCF adeptly adapts to various unseen images, accurately delineating changed objects—significantly outperforming the current state-of-the-art model. Moreover, as a zero-shot framework, GeSCF does not learn annotation biases in a dataset (see Fig. 5). For instance, GeSCF generates a more accurate and interpretable mask for the object and identifies unannotated changes, such as a sign, a basket, and a car. These results suggest that GeSCF’s outputs are more reasonable than baselines, even though its quantitative results are lower for some samples in the in-domain settings.
5.2 Ablation Studies
To understand the effectiveness of our contributions, we conducted comprehensive ablation studies across multiple datasets, including the standard SCD datasets and ChangeVPR (see Tab. 3). We utilized baselines trained on TSUNAMI [42], which demonstrated superior average performance over others.
Fine-tuning Strategy. We fine-tuned various SCD-based adapter networks on the frozen SAM image encoder, involving correlation layers [14], CHVA [11], and the feature merging module [50]. Our experiments reveal that fine-tuning the adaptor model on relatively small SCD datasets significantly undermines its zero-shot capabilities, resulting in a substantial performance drop compared to our proposed GeSCF. Notably, our initial pseudo-masks alone achieve an F1-score that surpasses the fine-tuned variants.
Effectiveness of the Proposed Modules. We validated the effectiveness of each proposed module, including the Initial Pseudo-mask Generation with adaptive threshold function, Geometric-Semantic Mask Matching with geometric intersection matching (GIM), and semantic similarity matching (SSM). The findings suggest that each introduced module positively influences GeSCF’s overall performance, with the initial pseudo-mask generation (adaptive) and GIM demonstrating significant enhancement.
Method Metric
F1-score mIoU
Uni-temporal Training
SAM [23] + Corr. layers [14] 28.7 19.7
SAM [23] + CHVA [11] 30.4 21.1
SAM [23] + MTF [50] + MSF [50] 40.1 30.2
Bi-temporal Training
SAM [23] + Corr. layers [14] 30.0 21.1
SAM [23] + CHVA [11] 31.0 22.0
SAM [23] + MTF [50] + MSF [50] 40.3 30.3
GeSCF (Ours)
Pseudo. only (fixed) 43.0 26.3
Pseudo. only (adaptive) 53.9 34.9
Pseudo. + GIM 65.0 47.6
Pseudo. + GIM + SSM (ALL) 65.9 48.8
Table 3: Ablation study results.
Method Metric
F1-score Precision Recall
ISFA [52] 32.9 29.8 36.8
DSFA [16] 33.0 24.2 51.9
DCAE [6] 33.4 35.7 31.4
OBCD [55] 34.3 29.6 40.7
KPCA-MNet [53] 36.7 29.5 48.5
DCVA [40] 36.8 29.6 48.7
GeSCF (Ours) 48.0 52.2 44.4
Table 4: Quantitative comparison with unsupervised remote sensing CD methods on the SECOND (test) benchmark.
5.3 Beyond Scene Change Detection
Although our research primarily focuses on natural scene CDs and each CD field tends to focus on its own [25], we discovered that GeSCF can also be applied to zero-shot remote sensing CD. Following the standard evaluation metrics of the remote sensing domain, we evaluated our model against several unsupervised methods [52, 16, 6, 55, 53, 40] on the SECOND [59] benchmark (see Tab. 4). Remarkably, our GeSCF outperforms other baselines with 48.0% in F1-score and a high precision of 52.2%—underscoring its zero-shot potential for different CD domains.
6 Conclusion
In this study, we attempted to address the generalization issue and temporal bias of scene change detection using the Segment Anything Model for the first time to the best of our knowledge; we define GeSCD along with the novel metrics, evaluation dataset (ChangeVPR), and an evaluation protocol to effectively assess the generalizability and temporal consistency of SCD models. Furthermore, we proposed GeSCF, a generalizable zero-shot approach for the SCD task using the localized semantics of the SAM that does not require costly SCD labeling. Our extensive experiments demonstrated that the proposed GeSCF significantly outperforms existing SCD models on standard SCD datasets and overwhelms them on the ChangeVPR dataset while achieving complete temporal consistency. We expect our methods can serve as a solid step towards robust and generalizable Scene Change Detection research.